please integrate the following
∫(ax+b)/(cx+d)2    I am unable to understandd its solution in book please explain properly 

Grade:11

1 Answers

Vikas TU
4 years ago
Dear student 
Writing this equatioon in partial fraction 
(ax + b )/(cx + d)2 = A / (cx + d)  + B / (cx + d)2
ax + b = A( cx + d) + B 
put x = -d/c 
-ad/c + b = (-ad + bc) /c 
put  x =0 
b = Ad + B 
b – B = b – (-ad + bc ) / c 
A = a/c 
Put the values and integrate yu will get :
a/clog|cx + d| – (-ad + bc )/c2 * 1/(cx + d)